### v1

@@ 2.3 Asymptotic bias of OLS
In the absence of strict exogeneity, the OLS estimator may suffer from asymptotic bias and, in some cases, may even be inconsistent.
@@
where $M_{\tilde{\ell}\ell}$ are entries of the $n\times n$ projection matrix $M$, and $Q=\lim\frac{1}{n}\sum_{\ell=1}^{n}\tilde{x}_{\ell}^{2}$.
This lemma is a relatively straightforward extension of the well-known Nickell bias (Nickell, 1981). Nickell’s original work focused on a more restricted setting commonly referred to as a “dynamic panel data” model.
@@
This is a special case of our Example 1 with the weak exogeneity assumption $\mathbb{E}[e_{it}\!\mid\!\{y_{is}:s<t\}]=0$. Consider the simplest case of a balanced panel with $T_{i}=T$ for all $i$. The projection matrix $M$ corresponds to demeaning within entities, and thus has a block-diagonal structure when observations are ordered by cluster. Under homoskedasticity with error variance $\sigma^{2}$, $\mathbb{E}[x_{\tilde{\ell}}e_{\ell}]=\mathbb{E}[y_{i,s-1}e_{it}]=\beta^{s-1-t}\sigma^{2}$ when $s>t$. Substituting these expressions into Lemma 1 yields the bias formula originally derived in Nickell (1981). It is well known that in this setting, OLS is inconsistent for fixed $T$, as the bias does not vanish. The leading term of the bias is $O(1/T)$, and if $T$ grows slower than the number of clusters $N$, the bias may dominate the variance asymptotically, resulting in invalid inference. $\Box$
The source of Nickell’s bias is well understood: differencing out fixed effects mixes observations from different periods, and future values of the regressor may correlate with current errors. Our framework is more general than the standard dynamic panel specification in (4). It is worth emphasizing that the result established by Nickell (1981)—and generalized in Lemma 1—is best understood as an asymptotic bias: a term which, when subtracted from the OLS estimator, yields a consistent (or asymptotically unbiased) estimator. Asymptotic bias is distinct from finite sample bias, as it ignores the randomness in the denominator of the OLS expression. While this randomness becomes negligible asymptotically, it is present in finite samples.
@@ 2.4 Design based modeling
Model (1) combined with Assumption 1 can be viewed as an instance of outcome modeling, since it specifies the outcome equation. In cross-sectional settings, however, OLS estimators perform well when either one of the outcome or treatment equations is correctly specified. The alternative approach—based on the treatment equation—is referred to as design-based and is commonly used when researchers understand the random assignment of treatment, such as in randomized controlled trials (RCTs). In this paper, we define a design-based model by specifying the following treatment equation, along with Assumption 2:
@@
Suppose a researcher wishes to estimate the effect of access to resources on individual outcomes by conducting an RCT in which individuals $\ell$ are randomly provided with additional resources. The researcher assumes that the treatment assignment equation is correctly specified as $x_{\ell}=w_{\ell}^{\prime}\delta_{x}+v_{\ell},$ where $p_{\ell}=w_{\ell}^{\prime}\delta_{x}$ denotes the propensity score, which may depend on individual characteristics as well as cluster fixed effects.
@@
where $F_{\ell}$ is a friend-selection matrix indicating which peers may influence individual $\ell$. The error term $e_{\ell}$ is assumed to be independent of all assignments $x$, but may be correlated within clusters to capture common shocks (e.g., village-level effects). In this case, the outcome modeling approach is complicated by the need to specify the functional form of the indirect effect. This effect might depend on whether any friends are treated, the number of treated friends, or the unobserved intensity of friendships. The quantity $g(W,F_{\ell}^{\prime}x)$ may vary systematically within clusters, and this variation may not be well captured by fixed effects or observed covariates. Consequently, if the indirect effect is absorbed into the unmodeled regression error term, the outcome model (1) may be misspecified.
@@
by the independence of treatment assignment. In this context, the exclusion restriction matrix $\mathcal{E}$ is the complement of the adjacency matrix of the friendship graph: it has zeros for direct friends and ones elsewhere. $\Box$
@@
The primary exposition in this paper is framed within the outcome modeling approach. However, all results can equivalently be reformulated under the design-based perspective with only minor modifications. We provide such restatements in the remarks following the main theorems. Additionally, where appropriate, we discuss the possibility of a doubly robust formulation, in which the researcher assumes that either the outcome equation or the treatment equation is correctly specified, without committing to which one.
@@ 3.1 Class of estimators
This subsection shows that no unbiased estimator exists in these settings with clustered data and suggests a somewhat weaker restriction that we call correct centering. Assume we have clustered data $\{y_{\ell},x_{\ell}\}$ generated according to model (1) under Assumption 1. The set of controls $W$, the cluster assignments $S_{i}$, and the exclusion restriction matrix $\mathcal{E}$ are given to the econometrician and treated as fixed. We consider the problem of estimating the parameter $\beta$. Let $\mathcal{F}$ denote the class of all joint distributions $F$ over $(x,y)$ that satisfy model (1) under Assumption 1.
@@
Consider an estimator $\hat{\beta}^{A}=\frac{x^{\prime}Ay}{x^{\prime}Ax},$ where $A$ is an $n\times n$ fixed matrix. [^1] Not every correctly centered estimator has the form $\hat{\beta}^{A}$, or is even linear in $y$. The independence between clusters imposes strong restrictions on the joint distribution of the data and allows higher-order $U$-statistics to be added to the numerator. For instance, if $(x_{i},y_{i}),\ i=1,\dots,4$ are observations from distinct (and thus independent) clusters, and $y_{i}=\beta x_{i}+e_{i}$, then $\mathbb{E}\big{[}x_{1}x_{2}y_{3}y_{4}-y_{1}y_{2}x_{3}x_{4}\big{]}=0.$ This observation connects to recent discussions in Hansen (2022) and Portnoy (2022); Lei and Wooldridge (2022). It is correctly centered if and only if $A$ satisfies the following conditions:
@@
If the data are generated from the design-based model (5) under Assumption 2, then the results of Lemma 4 continue to hold with a slightly modified (POP) condition. Specifically, we must replace (POP) with the following: (POP) $MA=A$. In the case of the doubly robust model, the corresponding condition becomes: (POP′′) $MA=AM=A$.
Numerous methods have been proposed to correct for Nickell bias in the standard dynamic panel model (4), or in its extensions with a small number of additional strictly exogenous regressors. See Okui (2021) and Bun and Sarafidis (2015) for recent surveys. Our approach aligns with previously proposed methods based on internal instruments—e.g., Anderson and Hsiao (1981), Arellano and Bond (1991), and Ahn and Schmidt (1995).
@@ 3.2 Asymptotic efficiency considerations
The function $V(A)$ represents the asymptotic variance of the IV estimator under strong identification, i.e., when ${x^{\prime}Ax}/{\mathbb{E}[x^{\prime}Ax]}\xrightarrow{p}1$. This efficiency criterion follows the classical approach of Chamberlain (1987). Lemma 5 shows that, under ideal conditions, our proposed estimator using $A^{*}$ from (8) achieves asymptotic efficiency within the class of correctly centered estimators. The two ideal conditions are: first, conditional homoskedasticity, $\mathbb{E}[ee^{\prime}\!\mid\!x]=\sigma^{2}I_{n}$, which implies that the optimal instrument $z^{*}$ minimizes the first stage prediction error over all $z=A^{\prime}x$ with $A\in\mathcal{A}$:
@@
It is natural that the efficiency of an estimator depends on the distributional properties of the regressor $x$, especially in settings like ours where $x$ must be treated as random. Any prior information—such as the scale of $x$ or its dynamic dependence—can, in principle, be used to improve efficiency, for example, through cluster-specific weighting or time transformations (as in GLS). However, caution is needed: the matrix $A$ must remain independent of the randomness in either $x$ or $e$ to avoid bias, such as the many-instrument bias observed in the two-step GMM estimator of Arellano and Bond (1991). If additional structure is available, such as $\mathbb{E}[xx^{\prime}]=\Omega$, then asymptotic efficiency is attained by solving:
@@ 3.3 Solution to the optimization problem
and estimates the regression $y_{it}^{*}=\beta x_{it}^{*}+e_{it}^{*}$ using $x_{it}$ as an instrument. This estimator lies within the class proposed by Arellano and Bond (1991), but applies more broadly. See also Hayashi and Sims (1983); Arellano and Bover (1995).
@@
and for $t=1$, we use the full mean: $y_{i1}^{*}=y_{i1}-\frac{1}{T_{i}}\sum_{s=1}^{T_{i}}y_{is},~x_{i1}^{*}=x_{i1}-\frac{1}{T_{i}}\sum_{s=1}^{T_{i}}x_{is}.$ We then estimate the regression $y_{it}^{*}=\beta x_{it}^{*}+e_{it}^{*}$ using $x_{it}$ as the instrument.
@@
In Example 6, we have $\operatorname{tr}(A_{ii})=T_{i}-1-\frac{1}{T_{i}}$; the term $\frac{1}{T_{i}}$ reflects the loss of information due to robustness against one-period feedback. In Example 5, $\operatorname{tr}(A_{ii})=T_{i}-1-\sum_{t=2}^{T_{i}}\frac{1}{t}$. With $\sum_{t=2}^{T_{i}}\frac{1}{t}\geq\log(T_{i}+1)-\log(2)$, this shows a substantially greater efficiency loss when guarding against dependence on any future regressor. Finally, in a panel model with fixed effects, imposing contemporaneous exogeneity only $\mathbb{E}[x_{\ell}e_{\ell}]=0$ leads to $A^{*}=0$. In this case, the model contains no usable identifying variation.
@@ 4.1 Importance of cross-dependence
where the term in brackets estimates $\gamma_{3}-\gamma_{1}$, and $(2,3)$ refers to the group of students who had teacher $g=2$ in year 1 and $g=3$ in year 2, and the upper bar denotes the average over the group. Similar transformations apply to $x_{i}^{*}$ and $e_{i}^{*}$. Our estimator $\hat{\beta}^{A^{*}}$ is IV on the equation $y_{i}^{*}=\beta x_{i}^{*}+e_{i}^{*}$, using $x_{i1}$ as an instrument: $\hat{\beta}^{A^{*}}-\beta=\frac{\sum_{i=1}^{N}x_{i1}e_{i}^{*}}{\sum_{i=1}^{N}x_{i1}x_{i}^{*}}$. But since $e_{i}^{*}$ depends on all $e_{jt}$, the terms $x_{i1}e_{i}^{*}$ are dependent across $i$. Here, $A^{*}$ is not block-diagonal, making standard inference approaches invalid and motivating the need for a different variance estimator. $\Box$
@@
Second, the OLS transformation for a student in group $(1,3)$ involves not only students with shared teachers ($g=1$ or $g=3$), but also those indirectly connected, e.g., students in $(2,4)$, who share a teacher with others in $(1,4)$ or $(2,3)$. More generally, we can view students as nodes in a network, where edges represent shared teachers. OLS demeaning propagates information across this network, creating strong dependencies within connected components and complicating standard inference methods. [^2] While OLS demeaning is efficient under the benchmark conditions of Lemma 5, one may prefer a less efficient estimator that induces less dependence. In Example 7, instead of using the full sample to estimate $\gamma_{3}-\gamma_{1}$, one could rely only on students with the same teacher history, e.g., $\overline{\Delta y}_{(1,3)}.$ This results in independent “super-clusters” defined by the teacher pairs: $(1,3)$, $(1,4)$, $(2,3)$, and $(2,4)$. One can go further by forming pairs of students with identical teacher histories and differencing them, creating many small, independent clusters, then using clustered standard errors on that level. However, this comes at a cost: the resulting estimator may be highly inefficient due to poor estimation of fixed effects. This strategy is explored in Verdier (2020).

### v2

@@ 2.3 Asymptotic bias of OLS
When strict exogeneity fails, the OLS estimator may suffer from asymptotic bias and, in some cases, may even be inconsistent.
@@
where $M_{\tilde{\ell}\ell}$ are entries of the $n\times n$ projection matrix $M$, and $Q=\operatorname{plim}_{n\rightarrow\infty}\frac{1}{n}\sum_{\ell=1}^{n}\tilde{x}_{\ell}^{2}$.
This lemma is a relatively straightforward extension of the well-known Nickell bias (nickell1981biases). Nickell’s original work focused on a more restricted setting commonly referred to as a “dynamic panel data” model.
@@
This is a special case of our Example 1 with the weak exogeneity restriction $\mathbb{E}[e_{it}\!\mid\!\{y_{is}:s<t\}]=0$ for all $i,t$. Consider the simplest case of a balanced panel with $T_{i}=T$ for all $i$. The projection matrix $M$ corresponds to demeaning within units, and thus has a block-diagonal structure when observations are ordered by cluster. Under homoskedasticity with error variance $\sigma^{2}$, $\mathbb{E}[x_{\tilde{\ell}}e_{\ell}]=\mathbb{E}[y_{is}e_{it}]=\beta^{s-t}\sigma^{2}$ when $s\geq t$. Substituting these expressions into Lemma 1 yields the bias formula originally derived in nickell1981biases. It is well known that in this setting, OLS is inconsistent for fixed $T$, as the bias does not vanish. The leading term of the bias is $O(1/T)$, and if $T$ grows slower than the number of clusters $N$, the bias may dominate the variance asymptotically, resulting in invalid inference. ∎
The source of Nickell’s bias is well understood: differencing out fixed effects mixes observations from different periods, and future values of the regressor may correlate with current errors. Our framework is more general than the standard dynamic panel specification in (4). It is worth emphasizing that the result established by nickell1981biases—and generalized in Lemma 1—is best understood as an asymptotic bias: a term which, when subtracted from the OLS estimator, yields a consistent (or asymptotically unbiased) estimator. Asymptotic bias is distinct from finite sample bias, as it ignores the randomness in the denominator of the OLS expression. While this randomness becomes negligible asymptotically, it is present in finite samples.
@@ 2.4 Design based modeling
Model (1) combined with Assumption 1 can be viewed as an instance of outcome modeling, since it specifies the outcome equation. In cross-sectional settings, however, OLS estimators perform well when either the outcome or the treatment equation is correctly specified. The alternative approach—based on the treatment equation—is referred to as design-based and is commonly used when researchers understand the random assignment of treatment, such as in randomized controlled trials (RCTs). In this paper, we define a design-based model by specifying the following treatment equation, along with Assumption 2:
@@
Suppose a researcher wishes to estimate the effect of access to resources on individual outcomes by conducting an RCT in which individuals $\ell$ are randomly provided with additional resources. The researcher assumes that the treatment assignment equation is correctly specified as $x_{\ell}=w_{\ell}^{\prime}\delta_{x}+v_{\ell},$ where $w_{\ell}^{\prime}\delta_{x}$ denotes the propensity score, which may depend on individual characteristics as well as cluster fixed effects.
@@
where $F_{\ell}$ encodes the friendship links relevant for individual $\ell$, so that $F_{\ell}^{\prime}x$ is the vector of treatment assignments for peers who may influence individual $\ell$. The error term $e_{\ell}$ is assumed to be independent of all assignments $x$, but may be correlated within clusters to capture common shocks (e.g., village-level effects). In this case, the outcome modeling approach is complicated by the need to specify the functional form of the indirect effect. This effect might depend on whether any friends are treated, the number of treated friends, or the unobserved intensity of friendships. The quantity $g(W,F_{\ell}^{\prime}x)$ may vary systematically within clusters, and this variation may not be well captured by fixed effects or observed covariates. Consequently, if the indirect effect is absorbed into the unmodeled regression error term, the outcome model (1) may be misspecified.
@@
by the independence of treatment assignment. In this context, the exclusion restriction matrix $\mathcal{E}$ is the complement of the adjacency matrix of the friendship graph: it has zeros for direct friends and ones elsewhere. ∎
@@
Figure 1: Bias of $\hat{\beta}^{\mathrm{LS}}$ in simulations
Notes. This figure presents the absolute bias of $\hat{\beta}^{\mathrm{LS}}$ (solid line) and its cluster-robust standard error (dotted line) in a design-based DGP with $n=500$ and varying spillover strength $\alpha$. The absolute bias of the proposed estimator $\hat{\beta}^{A^{*}}$ (described in Theorem 1, dashed line) is shown as a benchmark.
Figure 1 illustrates this asymptotic bias in a simulation designed to mimic the network-interference setting in Example 3. We consider a DGP based on a stochastic block model where clusters represent school classes and within-cluster edges represent friendships. The researcher observes the friendship network but not the intensity of each connection, denoted as weight $G_{\tilde{\ell}\ell}=G_{\ell\tilde{\ell}}\sim\mathrm{Exp}(1)$ for connected pairs and zero otherwise. Treatment is binary and assigned independently across units with probability $\mu_{i(\ell)}$, which varies across clusters. The outcome $y_{\ell}$ includes a spillover component depending on the treatment of friends weighted by unobserved friendship intensities: $y_{\ell}=\beta\,x_{\ell}+\alpha\sum_{\tilde{\ell}}G_{\ell\tilde{\ell}}\,x_{\tilde{\ell}}+e_{\ell}.$ The least-squares estimator controlling for cluster fixed effects exhibits substantial bias in this setting. Full details of the simulation design are provided in Appendix B. Despite random assignment of treatment, the OLS estimator can be highly misleading when interference is present. The bias arises from the correlation induced by spillovers within clusters, illustrating the Nickell-type bias in clustered regressions with network interference.
The simulation illustrates the usefulness of the design-based perspective in settings where the treatment assignment mechanism is credible but the outcome equation may be difficult to specify. The primary exposition in this paper is framed within the outcome modeling approach. However, all results can equivalently be reformulated under the design-based perspective with only minor modifications. We provide such restatements in the remarks following the main theorems. Additionally, where appropriate, we discuss the possibility of a doubly robust formulation, in which the researcher assumes that either the outcome equation or the treatment equation is correctly specified, without committing to which one.
@@ 3.1 Class of estimators
This subsection shows that unbiased estimation is generally impossible once regressors are random, even under standard orthogonality conditions, and motivates a weaker requirement that we call correct centering. Assume we have clustered data $\{y_{\ell},x_{\ell}\}$ generated according to model (1) under Assumption 1. The set of controls $W$, the cluster assignments $S_{i}$, and the exclusion restriction matrix $\mathcal{E}$ are given to the econometrician and treated as fixed. We consider the problem of estimating the parameter $\beta$. Let $\mathcal{F}$ denote the class of all joint distributions $F$ over $(x,y)$ that satisfy model (1) under Assumption 1.
@@
To better understand the concept, it is helpful to recall that in a standard cross-sectional overidentified IV regression, the two-stage least squares (2SLS) estimator is not correctly centered. When the number of instruments is large, the resulting bias can be of first-order importance and is known as the many-instrument problem. In contrast, a one-step just-identified IV estimator—as well as jackknife or leave-one-out IV estimators in overidentified cross-sectional settings—remains correctly centered whenever the instruments are exogenous.
@@
Consider an estimator $\hat{\beta}^{A}=\frac{x^{\prime}Ay}{x^{\prime}Ax},$ where $A$ is an $n\times n$ fixed matrix. [^1] Not every correctly centered estimator has the form $\hat{\beta}^{A}$, or is even linear in $y$. The independence between clusters imposes strong restrictions on the joint distribution of the data and allows higher-order $U$-statistics to be added to the numerator. For instance, if $(x_{i},y_{i}),\ i=1,\dots,4$ are observations from distinct (and thus independent) clusters, and $y_{i}=\beta x_{i}+e_{i}$, then $\mathbb{E}\big[x_{1}x_{2}y_{3}y_{4}-y_{1}y_{2}x_{3}x_{4}\big]=0.$ This observation connects to recent discussions in hansen2022modern and portnoy2022linearity; lei2022estimators. It is correctly centered if and only if $A$ satisfies the following conditions:
@@
If the data are generated from the design-based model (5) under Assumption 2, then the results of Lemma 4 continue to hold with a slightly modified (POP) condition. Specifically, we must replace (POP) with the following: (POP) $MA=A$. In the case of the doubly robust model, the corresponding condition becomes: (POP′′) $MA=AM=A$. ∎
Our approach is closely related to methods based on internal instruments for dynamic panel data, such as anderson1981estimation, arellano1991some, and ahn1995efficient. Both arellano1991some and ahn1995efficient stack all available exclusion restrictions and estimate the model using overidentified 2SLS or GMM. In contrast, we construct a just-identified linear combination of internal instruments and implement a one-step IV procedure. This delivers a correctly centered estimator. Overidentified procedures generally do not enjoy this property, as the combination of many moment conditions together with an estimated weighting matrix introduces additional randomness. When the degree of overidentification is large, this can result in substantial many-instrument bias (alvarez2003time). Different choices of the matrix $A$ satisfying restrictions (CC) and (POP) span a broad class of linear internal instruments and, in particular, include the infeasible optimally weighted instrument of arellano1991some as a special case.
@@ 3.2 Asymptotic efficiency considerations
The function $V(A)$ represents the asymptotic variance of the IV estimator under strong identification, i.e., when ${x^{\prime}Ax}/{\mathbb{E}[x^{\prime}Ax]}\xrightarrow{p}1$. This efficiency criterion follows the classical approach of chamberlain1987asymptotic. Lemma 5 shows that, under ideal conditions, our proposed estimator using $A^{*}$ from (8) achieves asymptotic efficiency within the class of correctly centered estimators. The two ideal conditions are: first, conditional homoskedasticity, $\mathbb{E}[ee^{\prime}\!\mid\!x]=\sigma^{2}I_{n}$, which implies that the optimal instrument $z^{*}$ minimizes the first stage prediction error over all $z=A^{\prime}x$ with $A\in\mathcal{A}$:
@@
It is natural that the efficiency of an estimator depends on the distributional properties of the regressor $x$, especially in settings like ours where $x$ must be treated as random. Any prior information—such as the scale of $x$ or its dynamic dependence—can, in principle, be used to improve efficiency, for example, through cluster-specific weighting or time transformations (as in GLS). However, caution is needed: the matrix $A$ must remain independent of the randomness in either $x$ or $e$ to avoid bias, such as the many-instrument bias observed in the two-step GMM estimator of arellano1991some. If additional structure is available, such as $\mathbb{E}[xx^{\prime}]=\Omega$, then asymptotic efficiency is attained by solving:
@@ 3.3 Solution to the optimization problem
and estimates the regression $y_{it}^{*}=\beta x_{it}^{*}+e_{it}^{*}$ using $x_{it}$ as an instrument. This estimator lies within the class proposed by arellano1991some, but applies more broadly. See also hayashi1983nearly; arellano1995another.∎
@@
and for $t=1$, we use the full mean: $y_{i1}^{*}=y_{i1}-\frac{1}{T_{i}}\sum_{s=1}^{T_{i}}y_{is},~x_{i1}^{*}=x_{i1}-\frac{1}{T_{i}}\sum_{s=1}^{T_{i}}x_{is}.$ We then estimate the regression $y_{it}^{*}=\beta x_{it}^{*}+e_{it}^{*}$ using $x_{it}$ as the instrument.∎
@@
In Example 6, we have $\operatorname{tr}(A_{ii})=T_{i}-1-\frac{1}{T_{i}}$; the term $\frac{1}{T_{i}}$ reflects the loss of information due to robustness against one-period feedback. In Example 5, $\operatorname{tr}(A_{ii})=T_{i}-1-\sum_{t=2}^{T_{i}}\frac{1}{t}$. With $\sum_{t=2}^{T_{i}}\frac{1}{t}\geq\log(T_{i}+1)-\log(2)$, this shows a substantially greater efficiency loss when guarding against dependence on any future regressor. Finally, in a panel model with fixed effects, assuming contemporaneous exogeneity only $\mathbb{E}[x_{\ell}e_{\ell}]=0$ leads to $A^{*}=0$. In this case, the model contains no usable identifying variation.
@@ 4.1 Importance of cross-dependence
where the term in brackets estimates $\gamma_{3}-\gamma_{1}$, and $(2,3)$ refers to the group of students who had teacher $g=2$ in year 1 and $g=3$ in year 2, and the upper bar denotes the average over the group. Similar transformations apply to $x_{i}^{*}$ and $e_{i}^{*}$. Our estimator $\hat{\beta}^{A^{*}}$ is IV on the equation $y_{i}^{*}=\beta x_{i}^{*}+e_{i}^{*}$, using $x_{i1}$ as an instrument: $\hat{\beta}^{A^{*}}-\beta=\frac{\sum_{i=1}^{N}x_{i1}e_{i}^{*}}{\sum_{i=1}^{N}x_{i1}x_{i}^{*}}$. But since $e_{i}^{*}$ depends on all $e_{jt}$, the terms $x_{i1}e_{i}^{*}$ are dependent across $i$. Here, $A^{*}$ is not block-diagonal, making standard inference approaches invalid and motivating the need for a different variance estimator. ∎
@@
Second, the OLS transformation for a student in group $(1,3)$ involves not only students with shared teachers ($g=1$ or $g=3$), but also those indirectly connected, e.g., students in $(2,4)$, who share a teacher with others in $(1,4)$ or $(2,3)$. More generally, we can view students as nodes in a network, where edges represent shared teachers. OLS demeaning propagates information across this network, creating strong dependencies within connected components and complicating standard inference methods. [^2] While OLS demeaning is efficient under the benchmark conditions of Lemma 5, one may prefer a less efficient estimator that induces less dependence. In Example 7, instead of using the full sample to estimate $\gamma_{3}-\gamma_{1}$, one could rely only on students with the same teacher history, e.g., $\overline{\Delta y}_{(1,3)}.$ This results in independent “super-clusters” defined by the teacher pairs: $(1,3)$, $(1,4)$, $(2,3)$, and $(2,4)$. One can go further by forming pairs of students with identical teacher histories and differencing them, creating many small, independent clusters, then using clustered standard errors on that level. However, this comes at a cost: the resulting estimator may be highly inefficient due to poor estimation of fixed effects. This strategy is explored in verdier2018estimation.